Pencils cost Rs.2 each, and notebooks cost Rs.3 each. Fatima has income of Rs.60. She is considering a consumption bundle containing pencils and 10 notebooks.

b. Pencils cost Rs.2 each, and notebooks cost Rs.5 each. Fatima has income of Rs.110. She is considering a consumption bundle containing 2() pencils and 10 notebooks.

Pencils cost Rs.3 each, and notebooks cost Rs.lO each. Fatima has income of Rs.50. She is considering a consumption bundle containing 10 pencils and 3 notebooks
1
Expert's answer
2021-08-09T11:04:31-0400

Solution:

a.). Budget constraint:

I = PxX + PyY

Where: I = Income = 60

            Px = Price of pencils = Rs.2

            Py = Price of notebooks = Rs.3

            X = Pencils

            Y = Notebooks

60 = 2X + 3Y

Notebooks (Y) = 10pcs

60 = 2X + 3(10)

60 = 2X + 30

60 – 30 = 2X

30 = 2X

X = 15 Pencils

Fatima’s consumption bundle will be = 15 pencils and 10 notebooks

 

b.). I = PxX + PyY

110 = 2X + 5Y

Pencils (X) = 2pcs

Notebooks (Y) = 10pcs

110 = 2(2) + 5(10)

110 = 4 + 50

110 > 54

Amount saved = 110 – 54 = Rs.56

Fatima will be able to save Rs.56 since the amount spent will be less than her consumption bundle. That is, her income is more than the costs of her consumption bundle.

 

c.). I = PxX + PyY

50 = 3X + 10Y

Pencils (X) = 10pcs

Notebooks (Y) = 3pcs

50 = 3(10) + 10(3)

50 = 30 + 30

50 < 60

Shortage = 60 – 50 = 10

Fatima will not be able to afford the consumption bundle since it costs more than the income she has.